# Computational bottlenecks for denoising diffusions

### v1

@@ Abstract
Denoising diffusions provide a general strategy to sample from a probability distribution $\mu$ in $\mathbb{R}^{d}$ by constructing a stochastic process $(\hat{\boldsymbol{x}}_{t}:t\geq 0)$ in ${\mathbb{R}}^{d}$ such that the distribution of $\hat{\boldsymbol{x}}_{T}$ at large times $T$ approximates $\mu$. The drift $\boldsymbol{m}:{\mathbb{R}}^{d}\times{\mathbb{R}}\to{\mathbb{R}}^{d}$ of this diffusion process is learned from data (samples from $\mu$) by minimizing the so-called score-matching objective. In order for the generating process to be efficient, it must be possible to evaluate (an approximation of) $\boldsymbol{m}(\boldsymbol{y},t)$ in polynomial time.
Is every probability distribution $\mu$, for which sampling is tractable, also amenable to sampling via diffusions? We provide evidence to the contrary by constructing a probability distribution $\mu$ for which sampling is easy, but the drift of the diffusion process is intractable —under a popular conjecture on information-computation gaps in statistical estimation. We further show that any polynomial-time computable drift can be modified in a way that changes minimally the score matching objective and yet results in incorrect sampling.
@@ 1.1 Background
Denoising diffusions [SE19, HJA20] have emerged as a central paradigm in generative artificial intelligence (AI). Given a target distribution $\mu$ on ${\mathbb{R}}^{d}$, we want to sample $\boldsymbol{x}\sim\mu$. Diffusions achieve this goal by generating trajectories of a stochastic process $(\hat{\boldsymbol{x}}_{t}:0\leq t\leq T)$ whose state $\hat{\boldsymbol{x}}_{T}$ at large $T$ is approximately distributed according to $\mu$. Crucially, the process $(\hat{\boldsymbol{x}}_{t}:0\leq t\leq T)$ is a diffusion whose coefficients can be estimated —at least in principle— from data $\boldsymbol{x}_{1},\dots,\boldsymbol{x}_{N}\sim\mu$ by solving a regression problem (empirical risk minimization with respect to square loss). Explicit knowledge of $\mu$ is not needed in order to generate a fresh sample $\boldsymbol{x}\sim\mu$ and powerful deep learning models can be deployed to approximate the coefficients of the diffusion. Even this brief description suggests a natural question:
Are there probability distributions $\mu$ for which generating the denoising diffusion is intrinsically harder than only generating samples from $\mu$?
This paper points at a positive answer. We construct a simple family of distributions $\mu$ for which: $(1)$ Sampling from $\mu$ is easy; $(2)$ Evaluating the coefficients of the diffusion is expected to be hard. More precisely, evaluation of these coefficients requires solving a statistical estimation problem for which a gap is expected to exist between optimal estimation error, and error achieved by polynomial time algorithms. We further show that any drift that is is nearly optima among polynomial time computable ones, can be modified in a way that leaves the score-matching objective nearly unchanged but results in incorrect sampling. Our construction is novel in this context and formalizes past observations based on statistical physics; see [MRTS07, RTS09, GDKZ24, HMP24] and Section 2.
We emphasize that this failure of denoising diffusions is computational of nature and is not driven by the difficulty of estimating the coefficients of the diffusion.
Denoising diffusions [^1] We follow the construction given in [Mon23], rather than the original one: the two differ by a change of variables. implement a discretization of the following stochastic differential equation (SDE), with initialization $\boldsymbol{y}_{0}=\boldsymbol{0}$
@@
where $(\boldsymbol{B}_{t})_{t\geq 0}$ is a standard Brownian motion and expectation in the second line is with respect to $\boldsymbol{x}\sim\mu$ independent of $\boldsymbol{g}\sim{\mathsf{N}}(\boldsymbol{0},\boldsymbol{I}_{d})$.
@@
A standard Euler discretization of the SDE (1) is given by (for $\hat{\boldsymbol{z}}_{t}\sim{\mathsf{N}}(\boldsymbol{0},\boldsymbol{I}_{d})$ independent of the past):
@@
with $\Delta$ a small stepsize. The algorithm outputs $\hat{\boldsymbol{x}}_{T}=\hat{\boldsymbol{m}}(\hat{\boldsymbol{y}}_{T},T)$ for some suitably large $T$. Here $\hat{\boldsymbol{m}}(\,\cdot\,,t)$ is the estimated drift. We will refer to this as ‘diffusion sampling.’
Diffusions reduces the problem of sampling to the problem of approximating the conditional expectation (2). The mapping $\boldsymbol{y}\mapsto\boldsymbol{m}(\boldsymbol{y},t)$ is the Bayes-optimal estimator of $\boldsymbol{x}$ in Gaussian noise:
@@
In words, we are given a Gaussian observation $\boldsymbol{y}_{t}\sim{\mathsf{N}}(t\boldsymbol{x},t{\boldsymbol{I}}_{d})$ (for a single $t$) and want to estimate $\boldsymbol{x}$ as to minimize mean square error (MSE). Here the minimization is over all measurable functions ${\boldsymbol{\varphi}}$.
In practice we are given a finite sample $(\boldsymbol{x}_{i})_{i\leq N}\sim_{iid}\mu$. We thus replace the expected error empirical error and the minimization over all functions, by minimization over a suitable function class:
minimize $\;\;\;\;\;\frac{1}{N}\sum_{i=1}^{N}\big{\|}{\boldsymbol{\varphi}}(\boldsymbol{y}_{i,t})-\boldsymbol{x}_{i}\big{\|}^{2}\,,$ (6)
subj. to $\;\;\;\;\;{\boldsymbol{\varphi}}\in\mathscrsfs{N}_{A}\,.$
where $\boldsymbol{y}_{i,t}=t\boldsymbol{x}_{i}+\sqrt{t}\boldsymbol{g}_{i}$ for i.i.d. standard Gaussian vectors $\boldsymbol{g}_{i}$.
In applications, $\mathscrsfs{N}_{A}$ is the set of neural networks with a certain architecture $A$ and (potentially) satifying suitable bounds on the norm of the weights. We will refer to the objective in Eqs. (5), (6) as to the score matching objective or the mean square error (MSE) objective. We emphasize that, for ${\boldsymbol{\varphi}}\in\mathscrsfs{N}_{A}$ the evaluation $\boldsymbol{y}\mapsto{\boldsymbol{\varphi}}(\boldsymbol{y})$ (a.k.a. ‘inference’ in deep learning jargon) can be implemented in polynomial time. This is indeed required in order to be able to implement the generative process (4).
@@ 1.2 Summary of results
As anticipated, we construct distributions in which sampling is easy but sampling via diffusions is hard, subject to conjectured hardness of certain statistical estimation problems. Indeed the construction is fairly simple. Let $\Omega_{n,k}:=\{\boldsymbol{u}\in\{0,\pm 1/\sqrt{k}\}^{n}|\|\boldsymbol{u}\|_{0}=k\}$ be the set of $0/\pm(1/\sqrt{k})$ unit vectors with $k$ nonzero entries (the notation $\|\boldsymbol{u}\|_{0}$ denotes the number of nonzeros in $\boldsymbol{u}$). We define the target distribution $\mu=\mu_{n,k}$ to be the distribution of
$\boldsymbol{x}=\boldsymbol{u}\boldsymbol{u}^{\sf T},\;\;\;\;\;\;\;\boldsymbol{u}\sim\operatorname{Unif}(\Omega_{n,k}).$ (7)
@@
Sampling from the target distribution $\mu$ is trivial. Just sample a vector with entries in $\{0,1/\sqrt{k},-1/\sqrt{k}\}$ and exactly $k$ non-zero entries, and let $\boldsymbol{x}=\boldsymbol{u}\boldsymbol{u}^{\sf T}$ as above. However, diffusion sampling requires to approximate the denoiser
@@
where $\boldsymbol{g}\sim{\mathsf{N}}({\boldsymbol{0}},{\boldsymbol{I}}_{n\times n})$. In other words, it requires to estimate a sparse rank-one matrix from a single observation in Gaussian noise. This is a well studied statistical estimation problem, closely related to the hidden clique problem, sparse PCA, and the Gaussian submatrix problem. Substantial rigorous evidence supports the claim that —for certain scalings of $k$, $t$ with $n$— polynomial-time algorithms cannot approach the Bayes-optimal error (for reference, see [BIS15, MW15, CLR17, BBH18, SW22].)
Our results suggest that sampling from $\mu_{n,k}$ using diffusions fails for $k\ll n$. More precisely, we will prove that there exists $\hat{\boldsymbol{m}}:{\mathbb{R}}^{n\times n}\times{\mathbb{R}}\to{\mathbb{R}}^{n\times n}$ such that:
  1. M1.
@@
In addition, we will prove a general reduction from estimation to diffusion sampling, establishing that if accurate diffusion sampling is possible in polynomial time, then near Bayes optimal estimation of $\boldsymbol{x}$ from $\boldsymbol{y}_{t}=t\boldsymbol{x}+\sqrt{t}\boldsymbol{g}$ must also be possible in polynomial time for all $t$. The contrapositive of this statement implies that if an information-computation gap exists, then diffusion sampling is impossible in polynomial time.
The rest of the paper is organized as follows. In Section 2 we motivate our setting and assumptions, and discuss some limitations of our results. In Section 3 we state formally our results (for technical reason we state two separate results depending on the growth of $k$ with $n$.) Section 4 presents the general reduction from estimation to diffusion sampling. Proofs are presented in Section 6 and the appendices.
Notations.
We write $\boldsymbol{W}\sim{\sf GOE}(n,v)$ if $\boldsymbol{W}=\boldsymbol{W}^{{\sf T}}$ is a random symmetric matrix with $(W_{ij})_{i\leq j\leq n}$ independent entries $W_{ii}\sim{\mathsf{N}}(0,2v)$, and $W_{ij}\sim{\mathsf{N}}(0,v)$ for $i<j$. For $v=1$, we use the shorter notation ${\sf GOE}(n)$. We say that $(\boldsymbol{W}_{t}:t\geq 0)$is a ${\sf GOE}(n)$ process if $\boldsymbol{W}_{t}\in{\mathbb{R}}^{n\times n}$ is a symmetric matrix with entries above and on the diagonal $(W_{t}(i,j):i<j\leq n;W_{t}(i,i)/\sqrt{2}:i\leq n;t\geq 0)$ forming a collection of $n(n+1)/2$ independent Brownian motions.
We use $C,C_{i},c_{i},\dots$ to denote absolute constants, whose value can change from line to line. We always consider sequences of sampling problem indexed by $n$, and will write (for instance) ‘there exists an estimator $\hat{\boldsymbol{m}}(\,\cdot\,)$’ instead of the more correct ‘there exists a sequence of estimators $\hat{\boldsymbol{m}}_{n}(\,\cdot\,)$ indexed by $n\in{\mathbb{N}}$.’
@@ 2 Discussion
Our results indicate that a standard application of denoising diffusions methodology will fail to sample from $\mu_{n,k}$ despite the sampling problem in itself is easy. On the other hand, the latent structure of the distribution $\mu_{n,k}$ can be exploited to construct a better algorithm. Namely, one can use diffusions to sample $\boldsymbol{u}\sim\operatorname{Unif}(\Omega_{n,k})$ (this is easy since the posterior expectation $\boldsymbol{m}(\boldsymbol{y},t)=\mathbb{E}[\boldsymbol{u}|t\boldsymbol{u}+\sqrt{t}\boldsymbol{g}=\boldsymbol{y}]$ is easy to compute) and then generate $\boldsymbol{x}=\boldsymbol{u}\boldsymbol{u}^{{\sf T}}$. However, identifying such latent structures from data is —in general— a hard problem, both from a statistical and from a computational viewpoint. As an example, one could consider a modification of the present problem whereby $\boldsymbol{u}$ is sparse in an unknown basis.
@@
We prove that there exists drifts $\hat{\boldsymbol{m}}(\,\cdot\,,t)$ that lead to poor generative sampling, despite being nearly optimal (among poly-time algorithms) in terms of the score matching objective (5). In particular, these bad drifts will be near optimal solutions of the empirical risk minimization problem (6), as long as $\mathscrsfs{N}_{A}$ is rich enough to approximate them. This happens because the class $\mathscrsfs{N}_{A}$ is a subset of the class of functions that can be evaluated in polynomial time.
We do not formally exclude the existence of drifts $\hat{\boldsymbol{m}}(\,\cdot\,,t)$ that also satisfy conditions M1 and M2 but yield good generative sampling. However if such drift existed, our results suggest that minimizing the score matching objective is not the right approach to find them.
@@
The subtlety in the previous point can be emphasized by the following remark. If we drop condition M2, i.e. we accept drifts that are bad from the point of view of the score matching, then it is possible to construct drifts that can be evaluated in polynomial time and yield good sampling. However, these are constructions that completely ignore the prescription (2). We provide one such construction of $\hat{\boldsymbol{m}}$ here, which guarantees that both $\hat{\boldsymbol{m}}(\hat{\boldsymbol{y}}_{t},t)$ and $\hat{\boldsymbol{y}}_{T}/T$ are approximate samples of $\boldsymbol{x}$ (the latter only for large $T$). Furthermore, this drift $\hat{\boldsymbol{m}}$ may depend on the noise stream $(\hat{\boldsymbol{z}}_{t})$ per Eq. (4), but relies on no additional source of randomness (the proof is presented in Appendix B.)
@@
Again, it is crucial to observe that the drift constructed in this proposition is artificial and has very poor value of the score-matching objective.
@@
Related work.
A number of groups proved positive results on diffusion sampling. Among others, [AMS22, CCL+23, MW23, LLT23, BDBDD23] provide reductions from diffusion sampling to score estimation, while [CLL23, SCK23, MW25, LHW24] establish end-to-end guarantees for classes of distributions $\mu$.
The computational bottleneck that we study here has been observed before in the context of certain Gibbs measures and Bayes posterior distributions [GDKZ24, AMS23, HMP24], and random constraint satisfaction problems [MRTS07, RTS09] (the latter papers use however sequential sampling rather than diffusion sampling).
Our work provides a simpler example and a rigorous formalization of the latter line of work. (Notice that despite superficial similarities, our construction is different from the examples of [GDKZ24, AMS23, HMP24], and allows us to build existing hardness results.)

### v2

@@ Abstract
Denoising diffusions sample from a probability distribution $\mu$ in $\mathbb{R}^{d}$ by constructing a stochastic process $(\hat{\boldsymbol{x}}_{t}:t\geq 0)$ in ${\mathbb{R}}^{d}$ such that $\hat{\boldsymbol{x}}_{0}$ is easy to sample, but the distribution of $\hat{\boldsymbol{x}}_{T}$ at large $T$ approximates $\mu$. The drift $\boldsymbol{m}:{\mathbb{R}}^{d}\times{\mathbb{R}}\to{\mathbb{R}}^{d}$ of this diffusion process is learned my minimizing a score-matching objective.
Is every probability distribution $\mu$, for which sampling is tractable, also amenable to sampling via diffusions? We provide evidence to the contrary by studying a probability distribution $\mu$ for which sampling is easy, but the drift of the diffusion process is intractable —under a popular conjecture on information-computation gaps in statistical estimation. We show that there exist drifts that are superpolynomially close to the optimum value (among polynomial time drifts) and yet yield samples with distribution that is very far from the target one.
@@ 1.1 Background
Diffusion sampling (DS) (Song and Ermon, 2019; Ho et al., 2020) has emerged as a central paradigm in generative artificial intelligence (AI). Given a target distribution $\mu$ on ${\mathbb{R}}^{d}$, we want to sample $\boldsymbol{x}\sim\mu$. Diffusions achieve this goal by generating trajectories of a stochastic process $(\hat{\boldsymbol{x}}_{t}:0\leq t\leq T)$ whose state $\hat{\boldsymbol{x}}_{T}$ at large $T$ is approximately distributed according to $\mu$. This suggests a natural question:
  • Q:
    Are there distributions $\mu$ for which sampling via diffusions fails even if sampling from $\mu$ is easy?
In order to explain how DS might fail, it is useful to recall the setup and introduce some notations [^1] We follow the formulation of Montanari (2023), which does not require time reversal.. The basic approach implements an approximation of the following stochastic differential equation (SDE), with initialization $\boldsymbol{y}_{0}=\boldsymbol{0}$
@@
where $(\boldsymbol{B}_{t})_{t\geq 0}$ is Brownian motion and in the second line $\boldsymbol{x}\sim\mu$ is independent of $\boldsymbol{g}\sim{\mathsf{N}}(\boldsymbol{0},\boldsymbol{I}_{d})$.
@@
A standard Euler discretization of the SDE (1) is given by (for $(\hat{\boldsymbol{z}}_{t})_{t\in{\mathbb{N}}\Delta}\sim_{iid}{\mathsf{N}}(\boldsymbol{0},\boldsymbol{I}_{d})$):
@@
with $\Delta$ a small stepsize. The algorithm outputs $\hat{\boldsymbol{x}}_{T}=\hat{\boldsymbol{m}}(\hat{\boldsymbol{y}}_{T},T)$ for some suitably large $T$. Here $\hat{\boldsymbol{m}}(\,\cdot\,,t)$ is the estimated drift.
Diffusions reduce the problem of sampling to the problem of approximating the conditional expectation (2). The mapping $\boldsymbol{y}\mapsto\boldsymbol{m}(\boldsymbol{y},t)$ is the Bayes-optimal estimator of $\boldsymbol{x}$ in Gaussian noise:
@@
In words, we are given a Gaussian observation $\boldsymbol{y}_{t}\sim{\mathsf{N}}(t\boldsymbol{x},t{\boldsymbol{I}}_{d})$ (for a single $t$) and want to estimate $\boldsymbol{x}$ as to minimize mean square error (MSE). This is also known as the ‘score-matching objective’. Minimization is over all measurable functions ${\boldsymbol{\varphi}}$.
In practice the minimization in Eq. (5) has to be modified for two reasons:
First, in general we do not know the distribution of $\boldsymbol{x}$ over which the expectation in Eq. (5) is taken, and instead we are given a finite sample $(\boldsymbol{x}_{i})_{i\leq N}\sim_{iid}\mu$. We thus replace the MSE by its empirical version:
minimize $\;\;\;\frac{1}{N}\sum_{i=1}^{N}\big{\|}{\boldsymbol{\varphi}}(\boldsymbol{y}_{i,t})-\boldsymbol{x}_{i}\big{\|}^{2}\,,\;\;\;\;\mbox{subj. to}\;\;{\boldsymbol{\varphi}}\in\mathscrsfs{N}\,,$ (6)
where $\boldsymbol{y}_{i,t}=t\boldsymbol{x}_{i}+\sqrt{t}\boldsymbol{g}_{i}$ for i.i.d. standard Gaussian vectors $\boldsymbol{g}_{i}$. Note that, for statistical reasons, the minimization over all functions must be replaced by by minimization over a suitable function class $\mathscrsfs{N}$ (e.g. neural nets). Changing the population problem (5) into the empirical risk minimization problem of Eq. (6) results into a statistical error that (ideally) vanishes as $N\to\infty$.
Second, the evaluation $\boldsymbol{y}\mapsto\hat{\boldsymbol{m}}(\boldsymbol{y},t)$ (a.k.a. ‘inference’ in deep learning jargon) must admit a polytime algorithm, in order to be able to implement the generative process (4). For this reason, $\mathscrsfs{N}$ must be a set of functions that can be evaluated in polynomial time. This is a purely computational constraint, and is present even if we have access to the population distribution (i.e., for $N=\infty$).
It is natural to suspect that the second constraint might lead to incorrect DS if near Bayes-optimal denoising cannot be achieved in polynomial time. The recent literature on ‘information-computation gaps’ provides a wealth of examples of statistical estimation problems for which evidence exists of a gap between Bayes optimal error and the optimal error achievable in polynomial time (Brennan et al., 2018; Bandeira et al., 2022; Celentano and Montanari, 2022; Schramm and Wein, 2022). Several of these problems take the form of estimating $\boldsymbol{x}\sim\mu$ from Gaussian observations $\boldsymbol{y}_{t}=t\boldsymbol{x}+\sqrt{t}\,\boldsymbol{g}$.
Koehler and Vuong (2024) already pointed out informally that denoising problems presenting an information-computation gap can result into a failure of DS. As a concrete example, they suggested the sparse spike Wigner model (see below). While this informal remark is natural, making it mathematically precise is far from obvious. In fact –strictly speaking– the remark is false. Roughly speaking, if sampling from $\mu$ is easy, then the drift $\boldsymbol{m}(\boldsymbol{y},t)$ can be constructed to return (for all $t\geq t_{0}$) a fixed random sample $\boldsymbol{x}\sim\mu$. Then the diffusion will sample correctly. However such $\boldsymbol{m}$ will be very far from an optimal denoiser. (See below for formal version of this argument.)
We also note that several earlier paper pointed at gaps between general sampling and DS or sequential sampling (Montanari et al., 2007; Ricci-Tersenghi and Semerjian, 2009; Ghio et al., 2024; Huang et al., 2024). None of these papers presented a formal claim either.
The present work fills this gap by providing mathematically rigorous versions of this claim in the setting of the sparse spiked Wigner model. We prove that there exists denoisers that are approximate optimizers (or sometimes optimizers) of the score matching objective (6) among polynomial time algorithms (up to an exponentially small error) and yet lead to completely incorrect sampling.
We emphasize that this failure of DS is computational of nature and purely related to the requirement to approximate the Bayes optimal denoiser $\boldsymbol{m}(\boldsymbol{y},t)$ by a polytime computable function.
@@ 1.2 Summary of results
The construction of the distribution $\mu$ is fairly simple. Let $B_{n,k}:=\{\boldsymbol{u}\in\{0,\pm 1/\sqrt{k}\}^{n}|\|\boldsymbol{u}\|_{0}=k\}$ be the set of $0/\pm(1/\sqrt{k})$ unit vectors with $k$ nonzero entries ($\|\boldsymbol{u}\|_{0}$ denotes the number of nonzeros in $\boldsymbol{u}$). We define the target distribution $\mu=\mu_{n,k}$ to be the distribution of
$\boldsymbol{x}=\boldsymbol{u}\boldsymbol{u}^{\sf T},\;\;\;\;\;\;\;\boldsymbol{u}\sim\operatorname{Unif}(B_{n,k}).$ (7)
@@
Sampling from the target distribution $\mu$ is trivial. Just sample a vector with entries in $\{0,1/\sqrt{k},-1/\sqrt{k}\}$ and exactly $k$ non-zero entries, and let $\boldsymbol{x}=\boldsymbol{u}\boldsymbol{u}^{\sf T}$. However, DS requires to approximate the denoiser
@@
where $\boldsymbol{g}\sim{\mathsf{N}}({\boldsymbol{0}},{\boldsymbol{I}}_{n\times n})$. In other words, it requires to estimate a sparse rank-one matrix from a single observation in Gaussian noise, sometimes referred to ‘sparse spiked Wigner,’ and closely related to the hidden clique problem or sparse PCA. Substantial rigorous evidence supports the claim that —for certain scalings of $k$, $t$ with $n$— polynomial-time algorithms cannot approach the Bayes-optimal error (Butucea et al., 2015; Ma and Wu, 2015; Cai et al., 2017; Brennan et al., 2018; Schramm and Wein, 2022).
We will prove three sets of results:
1. Near optimizers of score-matching can sample incorrectly. We prove that there exists $\hat{\boldsymbol{m}}:{\mathbb{R}}^{n\times n}\times{\mathbb{R}}\to{\mathbb{R}}^{n\times n}$ such that:
  • M1.
@@
2. Reduction from estimation to DS. We prove that if accurate DS is possible in polynomial time, then near Bayes optimal estimation of $\boldsymbol{x}$ from $\boldsymbol{y}_{t}=t\boldsymbol{x}+\sqrt{t}\boldsymbol{g}$ must also be possible in polynomial time for all $t$. The contrapositive of this statement implies that if an information-computation gap exists, then diffusion sampling is impossible in polynomial time.
3. All Lipschitz score-matching optimizers sample incorrectly. We prove that any denoiser that optimizes the score matching among polytime algorithms for $t\leq t_{1}$, and is $C/t$-Lipschitz for $t>t_{1}$ (for any constant $C$ a suitable $t_{1}$), samples incorrectly (under a modified version of $\mu_{n,k}$.)
The rest of the paper is organized as follows. In Section 2 we motivate our setting and assumptions, and discuss some limitations of our results. In Section 3 we state formally our results (for technical reason we state two separate results depending on the growth of $k$ with $n$.) Section 4 presents the general reduction from estimation to diffusion sampling. Section 5 proves that all Lipschitz score matching optimizers fail. A summary of notationss and proofs are presented in the appendices.
Throughout the paper, $C,C_{i},c_{i},\dots$ denote constants, whose value can change from line to line.
@@ 2 Discussion
Our results indicate that a standard application of denoising diffusions methodology will fail to sample from $\mu_{n,k}$ despite the sampling problem in itself is easy. On the other hand, the latent structure of the distribution $\mu_{n,k}$ can be exploited to construct a better algorithm. Namely, one can use diffusions to sample $\boldsymbol{u}\sim\operatorname{Unif}(B_{n,k})$ (this is easy since the posterior expectation $\boldsymbol{m}(\boldsymbol{y},t)=\mathbb{E}[\boldsymbol{u}|t\boldsymbol{u}+\sqrt{t}\boldsymbol{g}=\boldsymbol{y}]$ is easy to compute) and then generate $\boldsymbol{x}=\boldsymbol{u}\boldsymbol{u}^{{\sf T}}$. However, identifying such latent structures from data is (beyond the present example) a hard problem, both from a statistical and from a computational viewpoint.
@@
We prove that there exists drifts $\hat{\boldsymbol{m}}(\,\cdot\,,t)$ that lead to poor generative sampling, despite being nearly optimal (among poly-time algorithms) in terms of the score matching objective (5). In particular, these bad drifts will be near optimal solutions of the empirical risk minimization problem (6), as long as $\mathscrsfs{N}$ pnly contains polytime methods and is rich enough to approximate them. We further exclude the existence of Lipschitz drifts $\hat{\boldsymbol{m}}(\,\cdot\,,t)$ that also satisfy conditions M1 and M2 but yield good generative sampling.
In principle there could still be non-Lipschitz polytime drifts that are near score matching optimizers and sample well. However if such drift existed, our results suggest that minimizing the score matching objective is not the right approach to find them (since the difference in value with bad drifts will be superpolynomially small).
@@
As mentioned abovre, of we drop condition M2, i.e. we accept drifts that are bad from the point of view of the score matching, then it is possible to construct drifts that can be evaluated in polynomial time and yield good sampling. However, these are constructions that completely ignore the score-matching prescription (2). This is stated formally bnelow and proved in Appendix E.
@@
Again, the drift constructed in this proposition is artificial and has very poor value of the score-matching objective.
@@
Further related work.
A number of groups proved positive results on diffusion sampling. Alaoui et al. (2022); Chen et al. (2023b); Montanari and Wu (2023); Lee et al. (2023); Benton et al. (2023) provide reductions from diffusion sampling to score estimation. Chen et al. (2023a); Shah et al. (2023); Mei and Wu (2025); Li et al. (2024) give end-to-end guarantees for classes of distributions $\mu$.
The computational bottleneck that we study here has been observed before in the context of certain Gibbs measures and Bayes posterior distributions Ghio et al. (2024); Alaoui et al. (2023); Huang et al. (2024), and random constraint satisfaction problems Montanari et al. (2007); Ricci-Tersenghi and Semerjian (2009) (the latter papers use however sequential sampling rather than diffusion sampling).
Our work provides a simpler example and a rigorous formalization of the latter line of work. (Notice that despite superficial similarities, our construction is different from the examples of Ghio et al. (2024); Alaoui et al. (2023); Huang et al. (2024), and allows us to build existing hardness results.)